On December 1, Swifty Corporation has three DVD players left in stock. All are identical, all are priced to sell at $187. One of the three DVD players left in stock, with serial #1012, was purchased on June 1 at a cost of $104. Another, with serial #1045, was purchased on November 1 for $94. The last player, serial #1056, was purchased on November 30 for $84.

Calculate the cost of goods sold using the FIFO periodic inventory method assuming that two of the three players were sold by the end of December, Swifty Corporation year-end.

If Swifty Corporation used the specific identification method instead of the FIFO method, how might it alter its earnings by “selectively choosing” which particular players to sell to the two customers? What would Swifty’s cost of goods sold be if the company wished to minimize earnings? Maximize earnings?
Answer #1

Answer-

(a) Cost of goods sold (COGS) using FIFO

With FIFO, the earliest purchased item is sold out first, So,

COGS ($) = Purchase price of #1012 + Purchase price of #1045 = 104 + 94 = 198

(b) Earnings are minimized (maximized) when COGS is maximium (minimum).

(i) COGS is maximum when we choose the players #1012 and #1045.

COGS = $(104 + 94) = $198

(ii) COGS is minimum when we choose the players #1056 and #1045.

COGS = $(84 + 94) = $178